OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x + 3y = 1 3x - 4y = 10 A. 2 B. -1 C. 1 D. -2

OpenStudy (anonymous):

2x + 3y = 1 3x - 4y = 10 (-3)(2x + 3y = 1) (2)(3x - 4y = 10) -6x-9y=-3 6x-8y=20 -6x-9y+3 6x-8y-20 -17y=-17 y=-1 Plug y into any one of the equations above to find x

OpenStudy (anonymous):

the answer is b rite

OpenStudy (anonymous):

wait no i misunderstood

OpenStudy (anonymous):

the answer is A. :D

OpenStudy (anonymous):

Correct!